# Omnidirectional Multi-Object Tracking

### v1

@@ Abstract
Panoramic imagery, with its 360° field of view, offers comprehensive information to support Multi-Object Tracking (MOT) in capturing spatial and temporal relationships of surrounding objects. However, most MOT algorithms are tailored for pinhole images with limited views, impairing their effectiveness in panoramic settings. Additionally, panoramic image distortions, such as resolution loss, geometric deformation, and uneven lighting, hinder direct adaptation of existing MOT methods, leading to significant performance degradation. To address these challenges, we propose OmniTrack, an omnidirectional MOT framework that incorporates Tracklet Management to introduce temporal cues, FlexiTrack Instances for object localization and association, and the CircularStatE Module to alleviate image and geometric distortions. This integration enables tracking in large field-of-view scenarios, even under rapid sensor motion. To mitigate the lack of panoramic MOT datasets, we introduce the QuadTrack dataset—a comprehensive panoramic dataset collected by a quadruped robot, featuring diverse challenges such as wide fields of view, intense motion, and complex environments. Extensive experiments on the public JRDB dataset and the newly introduced QuadTrack benchmark demonstrate the state-of-the-art performance of the proposed framework. OmniTrack achieves a HOTA score of 26.92% on JRDB, representing an improvement of 3.43%, and further achieves 23.45% on QuadTrack, surpassing the baseline by 6.81%. The dataset and code will be made publicly available at https://github.com/xifen523/OmniTrack.
@@ 1 Introduction
To address these challenges, developing an MOT algorithm capable of comprehensive perception in panoramic images with large FoV is a pressing problem. To this end, this paper, for the first time, proposes an omnidirectional multi-object tracking framework, OmniTrack, specifically designed for such tasks in 360° panoramic imagery. OmniTrack unifies two mainstream MOT paradigms—Tracking-By-Detection (TBD) and End-To-End (E2E) tracking—and introduces a feedback mechanism that effectively reduces uncertainty in large FoV with rapid sensor motion, enabling fast and accurate target localization and association.
This framework consists of three core components: a CircularStatE Module, FlexiTrack Instance, and Tracklets Management. The CircularStatE Module is designed to mitigate wide-angle distortion and enhance consistency in lighting and color. The FlexiTrack Instance exploits the temporal continuity of objects, guiding the perception module to focus on key areas within the wide FoV and aiding in localization and association. This approach helps mitigate the difficulty of object localization in extra-wide FoV. The Tracklets Management module collects and manages trajectory data, providing prior knowledge to the FlexiTrack Instance. Through these components, OmniTrack unifies the two MOT paradigms: disabling data association within Tracklets Management results in an End-to-End tracker, OmniTrackE2E, while enabling association yields a TBD-style tracker, OmniTrackDA. By employing the same data association strategy, as shown in Fig. LABEL:fig:1 (c), the framework of OmniTrackDA achieves significantly stronger performance. Disabling both the FlexiTrack Instance and Tracklets Management reduces the system to a panoramic object detector, OmniTrackDet, as shown in Fig. LABEL:fig:1 (b).
Moreover, to support panoramic MOT research, we developed QuadTrack, a dataset collected with a $360^{\circ}{\times}70^{\circ}$ panoramic camera mounted on a quadrupedal robot. This mobile platform’s biomimetic gait introduces realistic, complex motion characteristics, challenging existing MOT algorithms. Collected across five campuses in two cities, QuadTrack includes $19,200$ images, encompassing a wide variety of dynamic, real-world scenarios. In contrast to typical MOT datasets [55, 8, 18, 5, 77, 15] that use static or linearly moving platforms, QuadTrack provides a new benchmark for evaluating MOT performance in large-FoV scenarios with rapid and non-linear sensor motion.
@@
  • To address the gap in omnidirectional multi-object tracking, we propose OmniTrack, a novel framework that unifies both E2E and TBD tracking paradigms. This approach reduces uncertainty and enhances perceptual and association performance in large-FoV scenarios.
  • We present QuadTrack, a new panoramic MOT dataset with complex motion dynamics, providing a challenging benchmark for wide-FoV multi-object tracking.
@@ 2 Related Work
Recently, researchers have recognized the advantages of omnidirectional images for tracking, particularly their ability to maintain continuous observation of targets without the out-of-view issues present in limited field-of-view setups. Jiang et al. [38] propose a $500$FPS omnidirectional tracking system using a three-axis active vision mechanism to capture fast-moving objects in complex environments. The 360VOT benchmark [33] is introduced for omnidirectional object tracking, focusing on spherical distortions and object localization challenges. Huang et al. [34] present 360Loc for omnidirectional localization that tackles cross-device challenges by generating lower-FoV query frames from 360° data. Another work by Xu et al. [73] introduces an extended bounding FoV (eBFoV) representation to alleviate spherical distortions in panoramic videos. Unlike previous methods, this work first explores extremely challenging large-FoV and intense-motion panoramic tracking for mobile robots, e.g., aiming to enhance the robot’s spatiotemporal understanding of objects in its surroundings.
Multi-object tracking. Object tracking primarily follows two paradigms: Tracking-By-Detection (TBD) [14, 22, 58, 83, 32, 50, 45, 57] and End-To-End (E2E) [19, 47, 25, 79]. Among these, TBD is currently one of the most prevalent, with frameworks following the design principles of SORT [71]. First, the detection network [27, 11] is used to locate bounding boxes for objects, then the target’s current position is predicted based on its historical trajectory, and the predicted results are associated with detection results [43]. Many subsequent works have refined this approach: DeepSORT [46] introduced a ReID model to incorporate appearance information for association, and ByteTrack [81] designed a confidence-based, stage-wise association strategy. Other methods [1, 76, 21] introduced motion compensation modules to mitigate camera motion, and OC-SORT [9] optimized the motion estimation module. Additionally, E2E methods have continued to evolve. TrackFormer [53] and MOTR [79] proposed transformer-based, End-to-End tracking approaches. Recent improvements [82, 50] have enhanced detector performance and improved data association accuracy in occlusion scenarios. Unlike existing methods that focus on narrow-FoV pinhole camera data with linear sensor motion, we address the challenges of MOT in wide-FoV scenarios, tackling issues such as geometric distortion and complex motion.
@@ 3 OmniTrack: Proposed Framework
In this section, we introduce OmniTrack, a panoramic multi-object tracking framework that addresses the unique challenges in wide-FoV images, including extensive search spaces, geometric distortion, resolution loss, and lighting inconsistencies. OmniTrack is designed with a feedback mechanism to iteratively refine object detection, integrating trajectory information back into the detector to enhance tracking stability across large-FoV scenes (Sec. 3.1). Specifically, to address these challenges, we propose three key components:
@@
  • CircularStatE Module (Sec. 3.4): Mitigates geometric distortion and improves consistency across the wide FoV, enhancing feature reliability.
@@ 3.1 Feedback Mechanism
The OmniTrack framework, illustrated in Fig. 1, differs from conventional MOT methods [81, 9, 21, 1], which separate object detection and data association. OmniTrack introduces a feedback mechanism that iteratively refines detections by integrating trajectory information back into the detector, enhancing stability for large-FoV tracking.
In traditional MOT, detection and association are decoupled, leading to higher entropy as each frame’s detection $H(x_{t})$ is calculated independently:
@@
where $H_{\text{feedback}}{<}H_{\text{independent}}$, indicating a reduction in uncertainty over time. This feedback-driven approach therefore enhances tracking stability in large-FoV scenarios.
Figure 2: The proposed CircularStatE Module fuses multi-scale features to generate learnable instances. The DynamicSSM Block mitigates distortions in large-FoV images, enhancing feature stability across uneven lighting and color distributions.
@@ 3.3 FlexiTrack Instance
As described in Equ. (2), the global association entropy is significantly high under wide-FoV conditions, making the association task challenging. Inspired by human tracking behavior, which focuses on probable regions based on past movement, we estimate the target’s current position by leveraging its historical trajectory. This approach avoids global search across the entire field of view, which is particularly beneficial in large-scale perception tasks. Based on this insight, we introduce FlexiTrack Instance.
Each FlexiTrack Instance shares the Decoder network structure with Learnable Instances, consisting of a feature vector $\mathcal{X}{\in}\mathbb{R}^{128}$ and an anchor $\mathcal{Y}{\in}\mathbb{R}^{128}$, as shown in Fig. 1. By sharing the decoder, FlexiTrack Instances can seamlessly adapt to various MOT paradigms, enhancing flexibility and allowing integration across different approaches without additional modifications. To enhance robustness, noise is added to both feature vectors and anchors during training, minimizing dependency on historical data and improving generalization:
@@
where $\mathcal{N}_{X}$ and $\mathcal{N}_{Y}$ represent the noise components added to the feature vector and anchor, respectively. To initialize all FlexiTrack Instance, let $\mathcal{I_{F}}$ denote the set of initial instance, and $N$ the total number of trajectories. Each instance $\mathcal{I_{F}}$ is composed of a feature vector $\mathcal{X}_{i}$ and an anchor $\mathcal{Y}_{i}$, as:
@@
Table 1: Typical datasets for 2D tracking. Abbreviations: [Uncaptioned image] (Autonomous Car), [Uncaptioned image] (Mobile Robot), [Uncaptioned image] (Quadruped Robot), [Uncaptioned image] (Internet images/videos), [Uncaptioned image] (Wheels), [Uncaptioned image] (Gait), [Uncaptioned image] (Stationary), Cov. (Coverage), Pano. (Panoramic image), Trk Len (Track Length), No. Seq (The number of sequences), No. Smp (The number of samples), and No. T (the number of tracks).
@@ 3.4 CircularStatE Module
The panoramic image provides an exceptionally wide FoV, capable of capturing 360° scenes. However, this inevitably introduces issues such as geometric distortions and inconsistencies in color and brightness in real-world high-dynamic-range scenes. To address these challenges, this paper proposes the CircularStatE Module, which alleviates distortions and improves the consistency of image features, thereby enhancing the performance of perception models.
@@ 4 QuadTrack: a Dynamic 360° MOT Dataset
Most existing MOT datasets [55, 18, 64] are captured using pinhole cameras, which are characterized by a narrow-FoV and linear sensor motion. However, when wide-FoV capture devices experience even slight movements, the entire scene can change drastically, posing significant challenges for object tracking. QuadTrack addresses this challenge by providing a benchmark specifically designed to test MOT algorithms under dynamic, non-linear motion conditions. It enables the evaluation of algorithm robustness in tracking objects with panoramic, non-uniform motion.
@@ 4.1 Dataset Collection and Challenges
To acquire a dataset with a wide FoV and complex motion dynamics, we utilized a quadruped robot dog as the data collection platform. This platform was selected for its biomimetic gait, which emulates the natural locomotion patterns of quadrupedal animals, introducing additional challenges for motion tracking due to its inherent complexity and variability. The robot measures $70cm{\times}31cm{\times}40cm$, with a maximum payload capacity of $7kg$. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we used a Panoramic Annular Lens (PAL) camera to capture wide-angle scenes with a FoV of $360^{\circ}{\times}70^{\circ}$. The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels, and supports a maximum output of $2448{\times}2048$ pixels at $40.5$FPS. Mounted on the quadruped robot (see Fig. 3 (b)), the camera ensures an unobstructed, optimal view. Using this platform, we conducted outdoor data collection across five campuses in two cities, capturing data in diverse environments.
With the biomimetic gait of the quadruped robot, the collected panoramic images naturally exhibited characteristic shaking, particularly along the Y-axis (Fig. 3 (c)). Compared to the JRDB dataset [51], our QuadTrack dataset introduces more complex motion challenges. Additionally, the data faces challenges such as uneven exposure, color inconsistencies due to the wide FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. More details can be found in the supplementary.
Figure 3: (a) shows the bounding box (bbox) size distribution for the training and validation sets, whereas (b) depicts the data collection platform and panoramic camera setup. (c) compares the Y-axis pixel positions of trajectories between the QuadTrack (Refer to caption) and JRDB [51] (Refer to caption) datasets, illustrating the significant vertical motion of the sensor in QuadTrack.
@@ 4.2 Data Distribution and Comparative Analysis
Unlike existing panoramic MOT datasets [55, 18, 28], which rely on pinhole cameras, QuadTrack, as shown in Tab. 1, is the first to be captured using a single $360^{\circ}$ panoramic camera. With a wider FoV ($360^{\circ}{\times}70^{\circ}$), QuadTrack significantly differs from traditional MOT datasets [55, 18]. In contrast to autonomous driving datasets [8, 77, 52], which often feature more predictable motion, QuadTrack incorporates complex, biologically inspired gait movements. Moreover, unlike internet-sourced datasets [64, 15], QuadTrack is designed to better reflect real-world application scenarios. While many existing datasets [54, 52, 8, 77] consist of short video sequences, QuadTrack emphasizes long-term tracking, with each video lasting $60$ seconds. To further challenge data association, we downsampled the dataset to $10$FPS, resulting in $600$ frames per sequence, spread across $32$ sequences. In total, QuadTrack includes $19,200$ frames and $189,876$ bounding boxes.
As illustrated in Fig. 3 (a), the distribution of both the training and test sets is consistent, ensuring a reliable and balanced evaluation of MOT methods. This similarity in the distribution between the sets reduces the potential for bias and allows for a more accurate comparison of model performance across varying conditions. The trajectories depicted in Fig. 3 (c) highlight the increased complexity of multi-object tracking under wide FoV conditions. Notably, the motion along the Y-axis is significantly more intense compared to JRDB, further complicating the tracking process.

### v2

@@ Abstract
Panoramic imagery, with its 360° field of view, offers comprehensive information to support Multi-Object Tracking (MOT) in capturing spatial and temporal relationships of surrounding objects. However, most MOT algorithms are tailored for pinhole images with limited views, impairing their effectiveness in panoramic settings. Additionally, panoramic image distortions, such as resolution loss, geometric deformation, and uneven lighting, hinder direct adaptation of existing MOT methods, leading to significant performance degradation. To address these challenges, we propose OmniTrack, an omnidirectional MOT framework that incorporates Tracklet Management to introduce temporal cues, FlexiTrack Instances for object localization and association, and the CircularStatE Module to alleviate image and geometric distortions. This integration enables tracking in panoramic field-of-view scenarios, even under rapid sensor motion. To mitigate the lack of panoramic MOT datasets, we introduce the QuadTrack dataset—a comprehensive panoramic dataset collected by a quadruped robot, featuring diverse challenges such as panoramic fields of view, intense motion, and complex environments. Extensive experiments on the public JRDB dataset and the newly introduced QuadTrack benchmark demonstrate the state-of-the-art performance of the proposed framework. OmniTrack achieves a HOTA score of 26.92% on JRDB, representing an improvement of 3.43%, and further achieves 23.45% on QuadTrack, surpassing the baseline by 6.81%. The established dataset and source code are available at https://github.com/xifen523/OmniTrack.
@@ 1 Introduction
To address these challenges, developing an MOT algorithm capable of comprehensive perception in panoramic images with panoramic FoV is a pressing problem. To this end, this paper, for the first time, proposes an omnidirectional multi-object tracking framework, OmniTrack, specifically designed for such tasks in 360° panoramic imagery. OmniTrack unifies two mainstream MOT paradigms—Tracking-By-Detection (TBD) and End-To-End (E2E) tracking—and introduces a feedback mechanism that effectively reduces uncertainty in panoramic FoV with rapid sensor motion, enabling fast and accurate target localization and association.
This framework consists of three core components: a CircularStatE Module, FlexiTrack Instance, and Tracklets Management. The CircularStatE Module is designed to mitigate wide-angle distortion and enhance consistency in lighting and color. The FlexiTrack Instance exploits the temporal continuity of objects, guiding the perception module to focus on key areas within the panoramic FoV and aiding in localization and association. This approach helps mitigate the difficulty of object localization in panoramic FoV. The Tracklets Management module collects and manages trajectory data, providing prior knowledge to the FlexiTrack Instance. Through these components, OmniTrack unifies the two MOT paradigms: disabling data association within Tracklets Management results in an End-to-End tracker, OmniTrackE2E, while enabling association yields a TBD-style tracker, OmniTrackDA. By employing the same data association strategy, as shown in Fig. LABEL:fig:1 (c), the framework of OmniTrackDA achieves significantly stronger performance. Disabling both the FlexiTrack Instance and Tracklets Management reduces the system to a panoramic object detector, OmniTrackDet, as shown in Fig. LABEL:fig:1 (b).
Moreover, to support panoramic MOT research, we developed QuadTrack, a dataset collected with a $360^{\circ}{\times}70^{\circ}$ panoramic camera mounted on a quadrupedal robot. This mobile platform’s biomimetic gait introduces realistic, complex motion characteristics, challenging existing MOT algorithms. Collected across five campuses in two cities, QuadTrack includes $19,200$ images, encompassing a wide variety of dynamic, real-world scenarios. In contrast to typical MOT datasets [55, 8, 18, 5, 77, 15] that use static or linearly moving platforms, QuadTrack provides a new benchmark for evaluating MOT performance in panoramic-FoV scenarios with rapid and non-linear sensor motion.
@@
  • To address the gap in omnidirectional multi-object tracking, we propose OmniTrack, a novel framework that unifies both E2E and TBD tracking paradigms. This approach reduces uncertainty and enhances perceptual and association performance in panoramic-FoV scenarios.
  • We present QuadTrack, a new panoramic MOT dataset with complex motion dynamics, providing a challenging benchmark for panoramic-FoV multi-object tracking.
@@ 2 Related Work
Recently, researchers have recognized the advantages of omnidirectional images for tracking, particularly their ability to maintain continuous observation of targets without the out-of-view issues present in limited field-of-view setups. Jiang et al. [38] propose a $500$FPS omnidirectional tracking system using a three-axis active vision mechanism to capture fast-moving objects in complex environments. The 360VOT benchmark [33] is introduced for omnidirectional object tracking, focusing on spherical distortions and object localization challenges. Huang et al. [34] present 360Loc for omnidirectional localization that tackles cross-device challenges by generating lower-FoV query frames from 360° data. Another work by Xu et al. [73] introduces an extended bounding FoV (eBFoV) representation to alleviate spherical distortions in panoramic videos. Unlike previous methods, this work first explores extremely challenging panoramic-FoV and intense-motion panoramic tracking for mobile robots, e.g., aiming to enhance the robot’s spatiotemporal understanding of objects in its surroundings.
Multi-object tracking. Object tracking primarily follows two paradigms: Tracking-By-Detection (TBD) [14, 22, 58, 83, 32, 50, 45, 57] and End-To-End (E2E) [19, 47, 25, 79]. Among these, TBD is currently one of the most prevalent, with frameworks following the design principles of SORT [71]. First, the detection network [27, 11] is used to locate bounding boxes for objects, then the target’s current position is predicted based on its historical trajectory, and the predicted results are associated with detection results [43]. Many subsequent works have refined this approach: DeepSORT [46] introduced a ReID model to incorporate appearance information for association, and ByteTrack [81] designed a confidence-based, stage-wise association strategy. Other methods [1, 76, 21] introduced motion compensation modules to mitigate camera motion, and OC-SORT [9] optimized the motion estimation module. Additionally, E2E methods have continued to evolve. TrackFormer [53] and MOTR [79] proposed transformer-based, End-to-End tracking approaches. Recent improvements [82, 50] have enhanced detector performance and improved data association accuracy in occlusion scenarios. Unlike existing methods that focus on narrow-FoV pinhole camera data with linear sensor motion, we address the challenges of MOT in panoramic-FoV scenarios, tackling issues such as geometric distortion and complex motion.
@@ 3 OmniTrack: Proposed Framework
In this section, we introduce OmniTrack, a panoramic multi-object tracking framework that addresses the unique challenges in panoramic-FoV images, including extensive search spaces, geometric distortion, resolution loss, and lighting inconsistencies. OmniTrack is designed with a feedback mechanism to iteratively refine object detection, integrating trajectory information back into the detector to enhance tracking stability across panoramic-FoV scenes (Sec. 3.1). Specifically, we propose the OmniTrack framework, which consists of three key components:
@@
  • CircularStatE Module (Sec. 3.4): Mitigates geometric distortion and improves consistency across the panoramic FoV, enhancing feature reliability.
@@ 3.1 Feedback Mechanism
The OmniTrack framework, illustrated in Fig. 1, incorporates a feedback mechanism that iteratively refines detections by integrating trajectory information back into the detector. This mechanism operates on the principle of reducing information entropy, thereby enhancing stability in Panoramic-FoV and improving MOT performance.
In traditional MOT [81, 9, 21, 1], detection and association are decoupled, leading to higher entropy as each frame’s detection $H(x_{t})$ is calculated independently:
@@
where $H_{\text{feedback}}{<}H_{\text{independent}}$, indicating a reduction in uncertainty over time. This feedback-driven approach thus enhances tracking stability in panoramic-FoV scenarios.
Figure 2: The proposed CircularStatE Module fuses multi-scale features to generate learnable instances. The DynamicSSM Block mitigates distortions in panoramic-FoV images, enhancing feature stability across uneven lighting and color distributions.
@@ 3.3 FlexiTrack Instance
As described in Eq. (2), the global association entropy is significantly high under panoramic-FoV conditions, making the association task challenging. Benefiting from the Feedback Mechanism (Sec. 3.1), which integrates trajectory information into the detector to reduce information entropy. This approach eliminates the need for global search across the entire field of view, making it especially effective for panoramic-scale perception tasks. Based on this insight, we introduce FlexiTrack Instance.
Each FlexiTrack Instance (see Fig. 1) shares the Decoder network structure with Learnable Instances, consisting of a feature vector $\mathcal{X}{\in}\mathbb{R}^{128}$ and an anchor $\mathcal{Y}{\in}\mathbb{R}^{128}$, as shown in Fig. 1. By sharing the decoder, FlexiTrack Instances can seamlessly adapt to various MOT paradigms, enhancing flexibility and allowing integration across different approaches without additional modifications. To enhance robustness, noise is added to both feature vectors and anchors during training, minimizing dependency on historical data and improving generalization:
@@
where $\mathcal{N}_{X}$ and $\mathcal{N}_{Y}$ represent the noise components added to the feature vector and anchor, respectively. To initialize all FlexiTrack Instances, let $\mathcal{I_{F}}$ denote the set of initial instances, and $N$ the total number of trajectories. Each instance $\mathcal{I_{F}}^{i}$ is composed of a feature vector $\mathcal{X}_{i}$ and an anchor $\mathcal{Y}_{i}$, as:
@@
Table 1: Typical datasets for 2D tracking. Abbreviations: [Uncaptioned image] (Autonomous Car), [Uncaptioned image] (Mobile Robot), [Uncaptioned image] (Quadruped Robot), [Uncaptioned image] (Internet images/videos), [Uncaptioned image] (Wheels), [Uncaptioned image] (Gait), [Uncaptioned image] (Stationary), Cov. (Coverage), Pano. (Panoramic camera), Trk Len (Track Length), No. Seq (The number of sequences), No. Smp (The number of samples), and No. T (the number of tracks).
@@ 3.4 CircularStatE Module
The panoramic image provides an exceptionally panoramic FoV, capable of capturing 360° scenes. However, this inevitably introduces issues such as geometric distortions and inconsistencies in color and brightness in real-world high-dynamic-range scenes. To address these challenges, this paper proposes the CircularStatE Module, which alleviates distortions and improves the consistency of image features, thereby enhancing the performance of perception models.
@@ 4 QuadTrack: a Dynamic 360° MOT Dataset
Most existing MOT datasets [55, 18, 64] are captured using pinhole cameras, which are characterized by a narrow-FoV and linear sensor motion. However, when panoramic-FoV capture devices experience even slight movements, the entire scene can change drastically, posing significant challenges for object tracking. QuadTrack addresses this challenge by providing a benchmark specifically designed to test MOT algorithms under dynamic, non-linear motion conditions. It enables evaluating algorithm robustness in tracking objects with panoramic, non-uniform motion.
@@ 4.1 Dataset Collection and Challenges
To acquire a dataset with a panoramic FoV and complex motion dynamics, we utilized a quadruped robot dog as the data collection platform. This platform was selected for its biomimetic gait, which emulates the natural locomotion patterns of quadrupedal animals, introducing additional challenges for motion tracking due to its inherent complexity and variability. The robot measures $70cm{\times}31cm{\times}40cm$, with a maximum payload capacity of $7kg$. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we used a Panoramic Annular Lens (PAL) camera to capture wide-angle scenes with a FoV of $360^{\circ}{\times}70^{\circ}$. The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels, and supports a maximum output of $2048{\times}2048$ pixels at $40.5$FPS. Mounted on the quadruped robot (see Fig. 3 (b)), the camera ensures an unobstructed, optimal view. Using this platform, the outdoor data collection spans morning, noon, afternoon, and evening, in diverse unconstrained environments across five campuses in two cities.
With the biomimetic gait of the quadruped robot, the collected panoramic images naturally exhibited characteristic shaking, particularly along the Y-axis (Fig. 3 (c) and (d)). Compared to the JRDB dataset [51], our QuadTrack dataset introduces more complex motion challenges. Additionally, the data faces challenges such as uneven exposure, color inconsistencies due to the panoramic FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. More details can be found in the supplementary.
Figure 3: (a) shows the bounding box (bbox) size distribution for the training and validation sets, whereas (b) depicts the data collection platform and panoramic camera setup. (c) and (d) compare the normalized Y-axis pixel positions of trajectories between the QuadTrack (Refer to caption) and JRDB [51] (Refer to caption) datasets, illustrating the significant vertical motion of the sensor in QuadTrack.
@@ 4.2 Data Distribution and Comparative Analysis
Unlike existing panoramic MOT datasets [55, 18, 28], which rely on pinhole cameras, QuadTrack, as shown in Tab. 1, is the first to be captured using a single $360^{\circ}$ panoramic camera. With a panoramic FoV ($360^{\circ}{\times}70^{\circ}$), QuadTrack significantly differs from traditional MOT datasets [55, 18]. In contrast to autonomous driving datasets [8, 77, 52], which often feature more predictable motion, QuadTrack incorporates complex, biologically inspired gait movements. Moreover, unlike internet-sourced datasets [64, 15], QuadTrack is designed to better reflect real-world application scenarios. While many existing datasets [54, 52, 8, 77] consist of short video sequences, QuadTrack emphasizes long-term tracking, with each video lasting $60$ seconds. To further challenge data association, we downsampled the dataset to $10$FPS, resulting in $600$ frames per sequence, spread across $32$ sequences. In total, QuadTrack includes $19,200$ frames and $189,876$ bounding boxes.
As illustrated in Fig. 3 (a), the distribution of both the training and test sets is consistent, ensuring a reliable and balanced evaluation of MOT methods. This similarity in the distribution between the sets reduces the potential for bias and allows for a more accurate comparison of model performance across varying conditions. The trajectories depicted in Fig. 3 (c) and (d) highlight the increased complexity of multi-object tracking under panoramic FoV conditions. Notably, the motion along the Y-axis is significantly more intense compared to JRDB [51], further increasing the difficulty of object detection and association.